OpenStudy (anonymous):

Medal Solve each system of equations by addition-subtraction, or by substitution. 3x + 2y = -4 -3x+y =7

Vocaloid (vocaloid):

hint: start by adding the two equations together

OpenStudy (anonymous):

-6x + 3y=3

OpenStudy (lynfran):

by adding the 2 equations note x will be eliminated

OpenStudy (lynfran):

3+(-3)=0

OpenStudy (anonymous):

so its 3y= 3

OpenStudy (anonymous):

y=1

OpenStudy (lynfran):

yes

OpenStudy (lynfran):

now sub y=1 into any of the above equations to find the value of x

OpenStudy (anonymous):

u get 3x + 2 = -4

OpenStudy (lynfran):

yea now take 2 frm both sides of the equation

OpenStudy (anonymous):

x= -2

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thank you so much

OpenStudy (lynfran):

welcome